At what angle, or slope in degrees will the hypotenus of this triangle ly?
The vertical leg measures 29"
The hypotnuse measures 63"
One angle of the trianlge is 90 degree

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

At what angle, or slope in degrees will the hypotenus of this triangle ly?
The vertical leg measures 29"
The hypotenuse measures 63"
One angle of the triangle is 90 degree
-------------
Draw the picture of the triangle.
slope = rise/run
------------------
run = sqrt(63^2-29^2) = 55.93 inches
----------------------
slope = 29/55.93 = 0.5185
=====================================
Cheers,
Stan H.

Answer by Alan3354(69443)

At what angle, or slope in degrees will the hypotenus of this triangle ly?
The vertical leg measures 29"
The hypotnuse measures 63"
One angle of the triangle is 90 degree
-------------
The other leg, the base is sqrt(63^2 - 29^2)
= sqrt(3128) =~ 55.9285
slope m = 29/55.9285 =~0.51852
------------
The angle is the arctan of m, =~ 27.407º
-----------------
Another method, without finding the slope, is
Angle = arcsin(29/63)
Angle = 27.407º